OpenStudy (freemap):

Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3 x = (7 – x) + 3 x – y = y + 3 x – (y + 3) = 7 (y + 3) – y = 7

OpenStudy (freemap):

My answer is A

OpenStudy (freemap):

I'am I right

OpenStudy (agent0smith):

No

OpenStudy (freemap):

why?

OpenStudy (agent0smith):

Idk how to answer that, other than, you didn't do it right. Show your work.

jimthompson5910 (jim_thompson5910):

notice the red x's \[\Large \color{red}{x} – y = 7 \] \[\Large \color{red}{x} =y+3\]

jimthompson5910 (jim_thompson5910):

substitution is another term for replacement example: a substitute teacher is a temporary replacement teacher. In mathematics if we had something like x+3 = 5 and we solved for x to get x = 2, then we can replace x with 2 to get 2+3 = 5. This is substitution in action

OpenStudy (freemap):

y+3 x-y=7

OpenStudy (freemap):

D

OpenStudy (agent0smith):

Yes.

OpenStudy (freemap):

Thank you

jimthompson5910 (jim_thompson5910):

yep you replace the first x with (y+3) since they are identical

OpenStudy (freemap):

Thank you